# AutoGUI: Scaling GUI Grounding with Automatic Functionality Annotations from LLMs

### v1

@@ 5.3 Failure Case Analysis
After analyzing the grounding failure cases, we identified several failure patterns in the fine-tuned models: a) difficulty in accurately locating small elements; b) challenges in distinguishing between similar but incorrect elements; and c) issues with recognizing icons that have uncommon shapes. Please refer to Sec. C.2 for details.
@@ 6 Conclusion
We propose AutoGUI, a scalable and automatic annotation pipeline aimed to produce massive UI element functionality annotations used to enhance UI understanding capabilities of open-source VLMs. The pipeline prompts an open-source LLM to generate element functionalities based on the UI content changes induced by interacting with the elements. To guarantee high quality, LLM-aided rejection and verification are introduced to remove invalid samples. Fine-tuned with the data collected by AutoGUI, the base models obtain strong UI grounding ability and exhibit data scaling effects. We hope that AutoGUI will open up possibilities for advancing the field of general UI agents.
@@ Appendix A Appendix
Section C: Additional experimental analysis including analysis of successful and failure cases on two benchmarks.
Section D and E: Limitations and Potential Societal Impact.
@@ A.1 Extra Statistics of the AutoGUI Dataset
A.2 Recording Interaction Trajectories on Web
@@
A.3 Recording Interaction Trajectories on Android Devices
@@
A.4 Functionality Annotation Details
@@
UI manipulation case We use a file-comparison library, DiffLib, to generate line-by-line differences of the AXtrees before and after interactions. To balance efficiency with annotation integrity, we limit the differences to 250 lines. In addition to the standard markings by DiffLib—addition, deletion, and unchanged status—we incorporate two additional change markers: ‘Repositioning’ and ‘Attribute Update’. These markers provide detailed information about UI content changes, essential for representing realistic structural variations. For example, without the attribute update marker, a clicked menu icon would erroneously appear as both deleted and added in the difference output, despite the menu element remaining in place. An example of this case is shown in Fig. D. The used prompt is shown in Tab. A.
@@
Table A: The functionality annotation prompt used in the AutoGUI pipeline in UI manipulation cases.
@@ 4. Your output MUST follow this format:
UI navigation case When an interacted element causes navigation to a new UI, the resultant changes are often extensive, potentially exceeding the context limit of an LLM and complicating the analysis of these changes. To handle this situation, UI description changes are used to predict functionalities. Concretely, the LLM is initially prompted to describe the UIs before and after interaction given UI AXTrees as input. Subsequently, the LLM then uses these descriptions to analyze content changes and predict the functionality of the interacted element. The description length of the AXTree is limited to 150 lines. An illustration of this process is shown in Fig. E. The corresponding prompt is detailed in Tab. B.
@@
Table B: The functionality annotation prompt used in the AutoGUI pipeline in UI navigation cases. This example shows how the LLM
@@ 5. Your output MUST follow this format:
A.5 Details of Rejecting Invalid Samples via Hand-Written Rules
@@
A.6 Details of Rejecting Invalid Samples via LLMs
To eliminate invalid samples before functionality annotation, the AutoGUI pipeline prompts the annotating LLM to also determine the validity of samples by analyzing the predictability of the UI content changes. The LLM evaluates each sample against three criteria: 1) Explicitness of Changes: This measures how clearly the changes indicate the element’s functionality. Changes that directly suggest functionality receive higher scores, while vague or irrelevant changes are not scored. 2) Relevance of Changes: This criterion assesses the significance of the modifications in relation to the element’s intended function. Highly related modifications obtain a high score. No scores for irrelevant or unrelated content changes. 3) Predictability of Outcome: This involves determining how anticipated the interaction outcome is based on the changes, considering common web conventions and user experience principles. Highly predictable changes obtain a high score, whereas moderate, unexpected, or counter-intuitive outcomes receive no score.
A.7 Details of LLM-Based Verification
To improve the quality of functionality annotations, the AutoGUI pipeline prompts two LLMs (i.e.g, Llama-3-70B and Mistral-7B-Instruct-v0.2) as verifiers to assign scores to samples based on how well the target elements adhere to their functionality annotations. The LLMs receive as the input a) the target element along with its surrounding UI content (up to 20 lines), b) the functionality annotation of this element, and c) the outcome of interacting with the element, either being the UI line-by-line differences (at most 250 lines) in manipulation cases or the UI description after the interaction in navigation cases. Given these inputs, the two LLMs generate two responses containing a score. Samples that do not achieve two full scores are discarded for higher quality of the AutoGUI dataset. The used prompt is shown in Tab. D and an example is illustrated in Fig. H.
Table C: The rejection prompt used in the AutoGUI pipeline in UI manipulation cases. This example shows how the LLM assigns a low score to a sample that exhibits meaningless and unpredictable UI content changes.
@@ 4. Your output MUST follow this format:
Table D: The self-verification prompt used in the AutoGUI pipeline in UI manipulation cases. This example shows how the LLM assigns a low score to the incorrect functionality.
@@
A.8 Details of Grounding/Captioning Task Generation
Table E: Examples of the task templates used by the AutoGUI pipeline to convert element-functionality pairs to functionality grounding/referring tasks. {functionality} and {point} are placeholders to be replaced with the real functionalities and target points.
Task Type Examples
Grounding For the element I describe in this page capture, can you predict their positions (with point)? {functionality}
In this web page image, please locate the element as I describe it (with point). {functionality}
Given a screenshot, I will describe a specific element; your task is to predict their locations (with point). {functionality}
Referring Describe the function of the element at {point} on the screen.
Detail the functionality of the UI element positioned at {point}.
What kind of input or interaction is expected at the point marked {point}?
After collecting the element-functionality pairs, the AutoGUI pipeline converts these pairs into functionality grounding and captioning tasks by formatting a multitude of task templates (several examples are shown in Tab. E). A functionality grounding task requires a VLM to output point coordinates of the element fulfilling the given functionality, while a captioning task demands that the VLM articulate a functionality description for an element, given its coordinates. It is important to note that each element-functionality pair is utilized to generate both a grounding task and a captioning task.
To optimize training efficiency and minimize token expenditure, all point coordinates are normalized within the range $[0,100)$. For tokenization, we employ the tokenizer from Qwen-VL-Chat without incorporating special tokens for the numerical range 0-99.
@@ B.1 Human Evaluation Details
To justify the efficacy of the AutoGUI pipeline, we conducted a comparative evaluation of annotation correctness between a trained human annotator and the AutoGUI system. The human annotator was a student proficient in using digital devices, ensuring familiarity with diverse user interfaces.
@@
During the training phase, we provided the annotator with detailed guidelines, including three specific criteria outlined in Fig 6, to ensure the clarity and correctness of their annotations. Additionally, we incorporated 15 invalid samples to instruct the human annotator on how to identify and exclude these cases during the evaluation process. These invalid samples encompassed scenarios such as incompletely loaded UIs, network failure incidents, login restrictions, and UIs displaying inappropriate content.
@@ B.2 Fine-Tuning Details
Qwen-VL-Chat Bai et al. (2023) and SliME Zhang et al. (2024b) are selected as the base models in the experiments. To investigate the scaling effects of our dataset, 25k, 125k, and the entirety of the 702k samples in the training split are used as training data in the three scaling experiments. For the first two smaller-scale experiments, a subset of the 702k data is randomly sampled.
@@
We train our UI-VLM based on the HuggingFace Transformers [^5] https://huggingface.co/docs/transformers/index and the PEFT library [^6] https://huggingface.co/docs/peft/index. The training configuration is shown in Tab. F and Tab. G.
C Additional Experimental Analysis
C.1 Growing Grounding Performance Brought by Scaling Data Size
@@
Figure I: Histograms of distances from predicted points to ground truth box centers. The distance from the normalized coordinate of a predicted point to its corresponding GT box center is calculated for all samples. Then, the histograms of these distances are illustrated to demonstrate the growing grounding performances brought by scaling the AutoGUI data size. The averaged distance for each experiment is displayed on the subplot title.
@@
To further investigate the benefit of scaling the AutoGUI functionality data, the histogram of distance from a predicted point to the ground truth box center is plotted for the 25k, 125k, and 702k experiments. The results in Fig. I demonstrate that the distance distributions become denser at lower ranges, suggesting that increasing the AutoGUI training data leads to consistently improved grounding performances.
C.2 Case Analysis on FuncPred Test Split
@@
Successful cases Fig. J demonstrates several examples of the grounding results from Qwen-VL trained with the 25k, 125k, and 702k AutoGUI data. The model trained with the 702k data (ours-702k) exhibits more accurate functionality grounding performance. For instance, Fig. J (a) shows that ours-702k predicts the point right on the target (The ‘Get an account’ button) while the other two models slightly miss the target. Case (c) shows that ours-702k correctly understands the functional intent to locate the WordPress logo, in contrast to the other models, which incorrectly focus on the text ‘Get WordPress’. Additionally, case (f) illustrates that ours-702k successfully locates the three-dot menu icon, aligning with the intent to expand a dropdown menu. These results suggest that increasing the AutoGUI training data enhances the model’s ability to understand complex functional intents and to recognize diverse iconic elements accurately.
@@
Failure cases To explore the limitations of our model, we analyze several failure cases across the scaling experiments, as shown in Fig. K. The primary failure cases comprise (1) Difficulty in accurately locating very small target elements, as illustrated by the tiny ‘Policy’ button in case (a); (2) Misunderstanding functional intents, as shown in case (b) where the three models fail to locate the element for account creation and case (g) where ours-702k mistakenly focuses on navigating to previous content instead of subsequent content; (3) Challenges in recognizing abstract iconic elements, as seen with the map style icon in case (d) and the compass icon in case (f).
@@
Figure J: Visualization of the successful functionality grounding examples for ours-625k. The ground truth bounding boxes, ours-625k predictions, ours-125k predictions, and ours-25k predictions are drawn in green, pink, blue, and orange, respectively.
Figure K: Visualization of failure examples in the scaling experiments. The ground truth bounding boxes, ours-625k predictions, ours-125k predictions, and ours-25k predictions are drawn in green, pink, blue, and orange, respectively.
C.3 Case Analysis on MoTIF Test Split
Figure L: Evaluation results of the model trained on 625k (blue dot) and 125k (red dot).
Figure M: Bad cases on MoTIF.
We evaluate the instruction following ability on MoTIF dataset. Our analysis focuses on two aspects: (1) what improvements our model can achieve with the scaling of our functionality dataset (Fig. L); and (2) in which scenarios our model still fails to achieve correct grounding (Fig. M).
Fig. L shows that the model can more accurately understand the action instruction and make meaningful localization as scaling improves from 125k to 702k. For instance, when the objective is to click sleep noise recording and click enable, the model can comprehend the semantics of this global objective and identify turn on. Additionally, the model can mitigate localization errors, such as the 702k being more accurately positioned on the target element (e.g., the icon of reservation) than the 125k. However, MoTIF still struggles with certain tasks. For example, as shown Fig. M, it has difficulty with localization in fine-grained steps for the instruction search for Kingston Drive and show me the route to it. It can be seen that the model does not effectively understand situations involving widget pop-ups (e.g., protocol and advertisement). This may be attributed to the weak semantic connection between pop-ups and the instruction. Furthermore, the model still falls short in precise localization. Enriching the dataset further could alleviate this issue.
Figure N: Comparing the annotations generated by a trained human annotator and the proposed AutoGUI pipeline. We can see that AutoGUI annotations are more detailed and clear than those by the human annotator.
D Limitations
AutoGUI Cannot Annotate UI Elements That Modify Content on the Internet To avoid causing potential contamination on the Internet and bearing unexpected responsibilities, we try our best to eliminate interaction samples that manipulate sensitive elements that probably modify contents on the Internet. For example, elements used to post comments, make purchases, and enter account information are discarded. Consequently, the AutoGUI pipeline mainly annotates elements that only support read-only functionalities.

### v2

@@ 4.3 Grounding Failure Case Analysis
After analyzing the grounding failure cases, we identified several failure patterns in the fine-tuned models: a) difficulty in accurately locating small elements; b) challenges in distinguishing between similar but incorrect elements; and c) issues with recognizing icons that have uncommon shapes. Please refer to Sec. D.2 for details.
5 Potential Use of AutoGUI Data
We apply our dataset to a downstream GUI agent task to demonstrate how our dataset can benefit GUI agents based on proprietary VLMs. The used benchmark is AITW (Rawles et al., 2023) which requires an agent to complete high-level user instructions on mobile apps. The metric is step accuracy, where a planned step is considered correct only if the action type and arguments match ground truths.
2-stage planning Following UGround (Gou et al., 2025), a planner model initially performs reasoning through task progress and UI content and then plans the next action. We prompt the planner to also describe the expected functionality of the target element for click actions. Next, a grounding model (Qwen2-VL-7B) trained with our functionality grounding tasks is used to locate the target according to the functionality description. See an example in Fig. 6 and additional details in Sec. C).
The results in Tab. 6 show that even the strong proprietary VLMs (e.g. Gemini) possess weak UI element grounding capability. Qwen2-VL trained with AutoGUI functionality grounding tasks can overtake the element grounding process of the proprietary models and help the planners achieve significantly higher step accuracy by correcting the target locations of the click actions.
Although this experiment is not designed to surpass expert models tailored for agent tasks, we hope it can facilitate further research of GUI agents with strong element grounding ability.
@@ 6 Conclusion
We propose AutoGUI, a scalable and automatic annotation pipeline aimed to produce massive UI element functionality annotations used to enhance UI grounding capabilities of VLMs. The pipeline prompts an open-source LLM to generate element functionalities based on the UI content changes induced by interacting with the elements. LLM-aided rejection and verification are used to guarantee high quality. Fine-tuned with the data collected by AutoGUI, the base models obtain stronger UI grounding ability and exhibit data scaling effects. We hope that AutoGUI will open up possibilities for advancing the field of UI agents.
Limitations
AutoGUI Cannot Annotate UI Elements That Modify Content on The Internet. To avoid causing potential contamination on the Internet and bearing unexpected responsibilities, we try our best to eliminate interaction samples that manipulate sensitive elements that probably modify contents on the Internet. For example, elements used to post comments, make purchases, and enter account information are discarded. Consequently, the AutoGUI pipeline mainly annotates elements that only support read-only functionalities.
Acknowledgments
This work was supported in part by the National Key R&D Program of China (No. 2022ZD0160102), the National Natural Science Foundation of China (No. U21B2042, No. 62320106010), and in part by the 2035 Innovation Program of CAS.
@@ Appendix A Appendix
Section C: Details of potential use of the AutoGUI dataset.
Section D: Additional experimental analysis including analysis of successful and failure cases on two benchmarks.
Section E: Limitations and Potential Societal Impact.
@@ A.1 Extra Statistics of the AutoGUI Dataset
The approach to avoidding overlap between train and test data: Since our focus is on annotating contextual functionality for GUI elements, we define two elements as distinct if they serve different functions within their respective contexts. For example, two "magnifier" buttons on the same GUI might have different roles—one for zooming in and the other for searching. To ensure no contamination, we investigated whether the test elements appeared in the training set by checking if bounding box overlapping occurred on the same GUIs. After this analysis, we found no such overlap.
A.2 License of The Used Artifacts
The licenses of the data sources on which the AutoGUI dataset is built are listed in Tab. A. These sources are all allowed to be used for academic research.
Table A: License or terms for use and/or distribution of the used artifacts in this work.
Artifacts License URLs containing Term-of-Use or other license information
Common Crawl CC BY https://commoncrawl.org/terms-of-use
AndroidControl Trajectories Apache 2.0 https://github.com/google-research/google-research/tree/41e4f1cbe1db648feb518a60501f638d9c8b25f2/android_control
Mobile Views Trajectories MIT https://huggingface.co/datasets/mllmTeam/MobileViews
A.3 Recording Interaction Trajectories on Web
@@
A.4 Recording Interaction Trajectories on Android Devices
@@
A.5 Functionality Annotation Details
@@
UI manipulation case We use a file-comparison library, DiffLib, to generate line-by-line differences of the AXtrees before and after interactions. To balance efficiency with annotation integrity, we limit the differences to 250 lines. In addition to the standard markings by DiffLib—addition, deletion, and unchanged status—we incorporate two additional change markers: ‘Repositioning’ and ‘Attribute Update’. These markers provide detailed information about UI content changes, essential for representing realistic structural variations. For example, without the attribute update marker, a clicked menu icon would erroneously appear as both deleted and added in the difference output, despite the menu element remaining in place. An example of this case is shown in Fig. D. The used prompt is shown in Tab. B.
@@
Table B: The functionality annotation prompt used in the AutoGUI pipeline in UI manipulation cases.
@@ 4. Your output MUST follow this format:
UI navigation case When an interacted element causes navigation to a new UI, the resultant changes are often extensive, potentially exceeding the context limit of an LLM and complicating the analysis of these changes. To handle this situation, UI description changes are used to predict functionalities. Concretely, the LLM is initially prompted to describe the UIs before and after interaction given UI AXTrees as input. Subsequently, the LLM then uses these descriptions to analyze content changes and predict the functionality of the interacted element. The description length of the AXTree is limited to 150 lines. An illustration of this process is shown in Fig. E. The corresponding prompt is detailed in Tab. C.
@@
Table C: The functionality annotation prompt used in the AutoGUI pipeline in UI navigation cases. This example shows how the LLM
@@ 5. Your output MUST follow this format:
A.6 Details of Rejecting Invalid Samples via Hand-Written Rules
@@
A.7 Details of Rejecting Invalid Samples via LLMs
Rejection process To eliminate invalid samples before functionality annotation, the AutoGUI pipeline prompts the annotating LLM to also determine the validity of samples by analyzing the predictability of the UI content changes. The LLM evaluates each sample against three criteria: 1) Explicitness of Changes: This measures how clearly the changes indicate the element’s functionality. Changes that directly suggest functionality receive higher scores, while vague or irrelevant changes are not scored. 2) Relevance of Changes: This criterion assesses the significance of the modifications in relation to the element’s intended function. Highly related modifications obtain a high score. No scores for irrelevant or unrelated content changes. 3) Predictability of Outcome: This involves determining how anticipated the interaction outcome is based on the changes, considering common web conventions and user experience principles. Highly predictable changes obtain a high score, whereas moderate, unexpected, or counter-intuitive outcomes receive no score.
A.8 Details of LLM-Based Verification
Verification process To improve the quality of functionality annotations, the AutoGUI pipeline prompts two LLMs (i.e.g, Llama-3-70B and Mistral-7B-Instruct-v0.2) as verifiers to assign scores to samples based on how well the target elements adhere to their functionality annotations. The LLMs receive as the input a) the target element along with its surrounding UI content (up to 20 lines), b) the functionality annotation of this element, and c) the outcome of interacting with the element, either being the UI line-by-line differences (at most 250 lines) in manipulation cases or the UI description after the interaction in navigation cases. Given these inputs, the two LLMs generate two responses containing a score. Samples that do not achieve two full scores are discarded for higher quality of the AutoGUI dataset. The used prompt is shown in Tab. E and an example is illustrated in Fig. H.
Table D: The rejection prompt used in the AutoGUI pipeline in UI manipulation cases. This example shows how the LLM assigns a low score to a sample that exhibits meaningless and unpredictable UI content changes.
@@ 4. Your output MUST follow this format:
Table E: The self-verification prompt used in the AutoGUI pipeline in UI manipulation cases. This example shows how the LLM assigns a low score to the incorrect functionality.
@@
A.9 Details of Grounding/Captioning Task Generation
After collecting the element-functionality pairs, the AutoGUI pipeline converts these pairs into functionality grounding and captioning tasks. A functionality grounding task requires a VLM to output point coordinates of the element fulfilling the given functionality, while a captioning task demands that the VLM articulate a functionality description for an element, given its coordinates. It is important to note that each element-functionality pair is utilized to generate both a grounding task and a captioning task.
@@ B.1 Human Evaluation Details
To justify the efficacy of the AutoGUI pipeline, we conducted a comparative evaluation of annotation correctness between a trained human annotator and the AutoGUI system. The human annotator was a graduate student proficient in using digital devices, ensuring familiarity with diverse user interfaces.
@@
During the training phase, we provided the annotator with detailed guidelines, including three specific criteria outlined in Fig I, to ensure the clarity and correctness of their annotations. Additionally, we incorporated 15 invalid samples to instruct the human annotator on how to identify and exclude these cases during the evaluation process. These invalid samples encompassed scenarios such as incompletely loaded UIs, network failure incidents, login restrictions, and UIs displaying inappropriate content.
@@
Figure I: The checking criteria used for comparing AutoGUI pipeline and the human annotator.
@@ B.2 Fine-Tuning Details
Qwen-VL-Chat Bai et al. (2023), SliME Zhang et al. (2024b), and Qwen2-VL-7B (Wang et al., 2024a) are selected as the base models in the experiments. To investigate the scaling effects of our dataset, 25k, 125k, and the entirety of the 702k samples in the training split are used as training data in the three scaling experiments. For the first two smaller-scale experiments, a subset of the 702k data is randomly sampled.
@@
Table H: The training hyper-parameters used for fine-tuning Qwen2-VL in the experiments.
Hyper-Parameter Value
Epoch 1
Global batch size 128
#GPUs 8
LoRA Rank 128
LoRA Alpha 256
Learning rate 3e-5
weight decay 0.0
ADAM Beta2 0.95
Warm-up ratio 0.03
LR scheduler Cosine
Model max length 2048
Frozen module ViT
DeepSpeed ZeRO-0
#Parameters
Trainable params: 322,961,408
All params: 8,614,337,024
Trainable%: 3.75
Data type BFloat16
@@
We train our models based on the HuggingFace Transformers [^5] https://huggingface.co/docs/transformers/index and the PEFT library [^6] https://huggingface.co/docs/peft/index. The training configurations are shown in Tab. F, Tab. G, and Tab. H.
Fine-tuning Qwen-VL-AutoGUI702k, SLiME-AutoGUI702k, Qwen2-VL-7B-AutoGUI702k, SeeClick w/ AutoGUI702k, UGround w/ AutoGUI702k consumed approximately 25 hours, 36 hours, 20 hours, 25 hours, 46 hours, respectively.
The framework used to fine-tune Qwen-VL and SeeClick w/ AutoGUI702k is the SeeClick codebase (Cheng et al., 2024); The framework used to fine-tune SLiME-AutoGUI702k is the SLiME codebase (Zhang et al., 2024b); The framework used to fine-tune Qwen2-VL-7B-AutoGUI702k and UGround w/ AutoGUI702k is LLaMA-Factory (Zheng et al., 2024).
B.3 Samples of Benchmarks
For clarity, the benchmarks’ samples are visualized in Fig. J.
Figure J: Samples of the UI grounding benchmarks used in the experiments
C Details for Potential Use of AutoGUI Dataset
We mainly conduct 2-stage planning on the AITW (Rawles et al., 2023) benchmark to assess the benefits of our AutoGUI data on downstream agent tasks.
As illustrated in Fig. 6, a planner is utilized to conduct reasoning and step prediction while a grounding model locates target elements for the actions that require targets (click, long-press, and hover). For other actions like swipe, back, home, and input-text, the grounding model is not involved. As this experiment requires the planner to describe the expected functionality of target elements, we use strong proprietary VLMs, such as GPT-4o-mini and Gemini-2.0 as the planners. Expert models, such as OS-ATLAS (Wu et al., 2024), are not used as they typically have lost general instruction following capability after large-scale fine-tuning.
The results in Tab. 6 show that Qwen2-VL-7B trained with our functionality grounding tasks can help the planners to more accurately locate target elements.
D Additional Experimental Analysis
D.1 Growing Grounding Performance Brought by Scaling Data Size
@@
Figure K: Histograms of distances from predicted points to ground truth box centers. The distance from the normalized coordinate of a predicted point to its corresponding GT box center is calculated for all samples. Then, the histograms of these distances are illustrated to demonstrate the growing grounding performances brought by scaling the AutoGUI data size. The averaged distance for each experiment is displayed on the subplot title.
@@
To further investigate the benefit of scaling the AutoGUI functionality data, the histogram of distance from a predicted point to the ground truth box center is plotted for the 25k, 125k, and 702k experiments. The results in Fig. K demonstrate that the distance distributions become denser at lower ranges, suggesting that increasing the AutoGUI training data leads to consistently improved grounding performances.
D.2 Case Analysis on FuncPred Test Split
@@
Successful cases Fig. L demonstrates several examples of the grounding results from Qwen-VL trained with the 25k, 125k, and 702k AutoGUI data. The model trained with the 702k data (ours-702k) exhibits more accurate functionality grounding performance. For instance, Fig. L (a) shows that ours-702k predicts the point right on the target (The ‘Get an account’ button) while the other two models slightly miss the target. Case (c) shows that ours-702k correctly understands the functional intent to locate the WordPress logo, in contrast to the other models, which incorrectly focus on the text ‘Get WordPress’. Additionally, case (f) illustrates that ours-702k successfully locates the three-dot menu icon, aligning with the intent to expand a dropdown menu. These results suggest that increasing the AutoGUI training data enhances the model’s ability to understand complex functional intents and to recognize diverse iconic elements accurately.
@@
Failure cases To explore the limitations of our model, we analyze several failure cases across the scaling experiments, as shown in Fig. M. The primary failure cases comprise (1) Difficulty in accurately locating very small target elements, as illustrated by the tiny ‘Policy’ button in case (a); (2) Misunderstanding functional intents, as shown in case (b) where the three models fail to locate the element for account creation and case (g) where ours-702k mistakenly focuses on navigating to previous content instead of subsequent content; (3) Challenges in recognizing abstract iconic elements, as seen with the map style icon in case (d) and the compass icon in case (f).
@@
Figure L: Visualization of the successful functionality grounding examples for ours-625k. The ground truth bounding boxes, ours-625k predictions, ours-125k predictions, and ours-25k predictions are drawn in green, pink, blue, and orange, respectively.
Figure M: Visualization of failure examples in the scaling experiments. The ground truth bounding boxes, ours-625k predictions, ours-125k predictions, and ours-25k predictions are drawn in green, pink, blue, and orange, respectively.
D.3 Case Analysis on MoTIF Test Split
Figure N: Evaluation results of the model trained on 625k (blue dot) and 125k (red dot).
Figure O: Bad cases on MoTIF.
We evaluate the instruction following ability on MoTIF dataset. Our analysis focuses on two aspects: (1) what improvements our model can achieve with the scaling of our functionality dataset (Fig. N); and (2) in which scenarios our model still fails to achieve correct grounding (Fig. O).
Fig. N shows that the model can more accurately understand the action instruction and make meaningful localization as scaling improves from 125k to 702k. For instance, when the objective is to click sleep noise recording and click enable, the model can comprehend the semantics of this global objective and identify turn on. Additionally, the model can mitigate localization errors, such as the 702k being more accurately positioned on the target element (e.g., the icon of reservation) than the 125k. However, MoTIF still struggles with certain tasks. For example, as shown Fig. O, it has difficulty with localization in fine-grained steps for the instruction search for Kingston Drive and show me the route to it. It can be seen that the model does not effectively understand situations involving widget pop-ups (e.g., protocol and advertisement). This may be attributed to the weak semantic connection between pop-ups and the instruction. Furthermore, the model still falls short in precise localization. Enriching the dataset further could alleviate this issue.
Figure P: Comparing the annotations generated by a trained human annotator and the proposed AutoGUI pipeline. We can see that AutoGUI annotations are more detailed and clear than those by the human annotator.